# Robust Quantum Control using Reinforcement Learning from Demonstration

### v1

@@ I Introduction
Reinforcement learning (RL), a sub-field of machine learning, has been proposed to address the problems discussed above. [17, 18] . RL has been used to solve complex control problems  [19, 20, 21]. Reinforcement learning can transform control problems into optimization problems, making it particularly suitable for enhancing the fidelity of quantum control. Reinforcement learning utilizes direct interactions with quantum systems, thereby having the ability to avoid model bias. Model-free RL has been widely used in quantum control [22, 23, 24, 25]. However, model-free RL methods need to interact intensively with quantum systems to collect data to learn a good control sequence, which will impede the generalization of reinforcement learning methods to larger and more complex tasks. Another main difficulty for RL in solving quantum control problems is the high dimensionality of the control field action space. For a quantum control task, there are always multiple control Hamiltonians, and each Hamiltonian requires over hundreds of parameters to describe its pulse characteristics [26]. This result is in a high-dimensional parameter space. Random exploration of these high-dimensional spaces is exceptionally challenging, especially when learning from scratch. Model-based RL attempts to learn a system model and then leverages this model to generate additional data, thereby reducing the demand for samples from real quantum systems [27]. Whereas, as the dimension of quantum systems increases, learning a precise model of such systems tends to become a very challenging task.
@@ II.1 Reinforcement learning and pulse generation
Demonstration here means state-action pair $(s,a)$ generated by pre-trained models or human experts. Physical calibrated gates can also be used as demonstrations. Also, one can train a model by optimizing gate parameters rather than pulse shapes. This approach of training gate parameters has fewer parameters, and it’s easier for an RL agent to learn from scratch. Pulse shapes can also be generated using that pre-trained model. Pulses generated can be fed into an RL agent which optimizes pulse shapes to achieve higher fidelity, as shown in Fig.1(a), (b). RL methods can also be integrated with simulations. Although accessing the quantum state during control is not feasible in experiments. In simulations, the quantum state can be utilized to aid the training process. This pre-trained RL agent can also be used in experimental quantum control tasks.
@@ III.2 Complex state preparation of bosonic system
Then we use RLfD method to prepare a superposition of Fock states or coherent states. Currently, quantum error correction (QEC) is being extensively studied in such states [40, ni2023beating]. Here, we use PLfD method to generate a binomial code state given by $(|0\rangle+|4\rangle)/{\sqrt{2}}$ and a even cat stae $|\alpha\rangle+|-\alpha\rangle$ with $\alpha=2$. We consider a transmon qubit and a cavity in our physical model, which are dispersively coupled. With the consideration of high level leakage, qubit modes are considered as bosonic modes with kerr factor. The dimension of the Hilbert space is truncated to 3 and 7 for the qubit and cavity, respectively. We use a Hamiltonian with consideration of Kerr factors [41],
@@ IV CONCLUSIONS
Our analysis highlights the distinct features and advantages of both off-policy and on-policy reinforcement learning algorithms within the context of quantum control. Specifically, the off-policy algorithm, soft actor critic (SAC), exhibits higher sample efficiency. It requires fewer interactions with the environment to reach a specified performance threshold and is capable of learning from scratch. This superiority is attributed to the more informed and reasonable initial values provided by the demonstrations, which guide the agent toward more promising regions of the action space.
Conversely, the on-policy algorithm, proximal policy optimization (PPO), encounters difficulties in converging when learning from scratch due to its exploration strategy and reliance on on-policy data. Without demonstrations, PPO struggles to handle quantum control problems with high-dimensional action spaces, as evidenced by its stagnant average reward during training. Nevertheless, when equipped with demonstrations and a pre-trained policy network, PPO excels. The prior knowledge significantly enhances its practicality, allowing it to converge to high-performance policies. The integration of demonstrations ensures greater training stability and convergence. This feature is crucial, especially in complex environments where prior knowledge can be extremely beneficial. Our findings suggest that PPO, when combined with RLfD, can be extended to more complex quantum control problems with high-dimensional action space.
@@ A.2 PPO algorithm
The algorithm PPO with pre-train for Quantum Control is shown below.
@@
$H(\cdot|s_{t})$ is the entropy of $\pi_{\theta}(\cdot|s_{t})$, it’s used to maintain randomness of $\pi_{\theta}(\cdot|s_{t})$ to preserve the exploration ability. Entropy here is used to prevent the policy’s variance from being zero. Here $\lambda_{\rm{ent}}$ is the coefficient of entropy that controls the extent of exploration by influencing how much randomness is introduced into the policy. We finally show the comparison of pulses shapes generated by RL agent and original demonstrations in Fig 5. We can see that the pulse shapes generated by the RL agent are only slightly different from the original demonstrations. These subtle differences indicate that the agent is capable of fine-tuning the pulse shapes to better adapt to the environment. By making minor adjustments, the RL agent can compensate for system imperfections or calibration errors, leading to improved performance in the quantum control task.

### v2

@@ I Introduction
Reinforcement learning (RL), a sub-field of machine learning, has been proposed to address the problems discussed above. [17, 18] . RL has been used to solve complex control problems  [19, 20, 21]. Reinforcement learning can transform control problems into optimization problems, making it particularly suitable for enhancing the fidelity of quantum control. Reinforcement learning utilizes direct interactions with quantum systems, thereby having the ability to avoid model bias. Model-free RL has been widely used in quantum control and related tasks [22, 23, 24, 25, 26, 27]. However, model-free RL methods need to interact intensively with quantum systems to collect data to learn a good control sequence, which will impede the generalization of reinforcement learning methods to larger and more complex tasks. Another main difficulty for RL in solving quantum control problems is the high dimensionality of the control field action space. For a quantum control task, there are always multiple control Hamiltonians, and each Hamiltonian requires over hundreds of parameters to describe its pulse characteristics [28]. This result is in a high-dimensional parameter space. Random exploration of these high-dimensional spaces is exceptionally challenging, especially when learning from scratch. Model-based RL attempts to learn a system model and then leverages this model to generate additional data, thereby reducing the demand for samples from real quantum systems [29]. Whereas, as the dimension of quantum systems increases, learning a precise model of such systems tends to become a very challenging task.
@@ II.1 Reinforcement learning and pulse generation
Demonstration here means state-action pair $(s,a)$ generated by pre-trained models or human experts. Physical calibrated gates can also be used as demonstrations. Also, one can train a model by optimizing gate parameters rather than pulse shapes. This approach of training gate parameters has fewer parameters, and it’s easier for an RL agent to learn from scratch. Pulse shapes can also be generated using that pre-trained model. Pulses generated can be fed into an RL agent which optimizes pulse shapes to achieve higher fidelity, as shown in Fig. 1(a) and (b). RL methods can also be integrated with simulations. Although accessing the quantum state during control is not feasible in experiments. In simulations, the quantum state can be utilized to aid the training process. This pre-trained RL agent can also be used in experimental quantum control tasks.
@@ III.2 Complex state preparation of bosonic system
Then we use RLfD method to prepare a superposition of Fock states or coherent states. Currently, quantum error correction (QEC) is being extensively studied in such states [42, 43]. Here, we use PLfD method to generate a binomial code state given by $(|0\rangle+|4\rangle)/{\sqrt{2}}$ and a even cat stae $|\alpha\rangle+|-\alpha\rangle$ with $\alpha=2$. We consider a transmon qubit and a cavity in our physical model, which are dispersively coupled. With the consideration of high level leakage, qubit modes are considered as bosonic modes with kerr factor. The dimension of the Hilbert space is truncated to 3 and 7 for the qubit and cavity, respectively. We use a Hamiltonian with consideration of Kerr factors [44],
@@
In Fig. 3(d), we illustrate the influence of channel bandwidth on state preparation. Our results indicate that when the bandwidth is sufficiently wide, the final fidelity remains high. However, when the bandwidth becomes overly constrained, the final fidelity reduces significantly. Notably, the RLfD method can mitigate this degradation by enhancing the final fidelity, although its performance does not fully match that achieved without bandwidth limitations.
@@ IV CONCLUSIONS
Our analysis highlights the distinct features and advantages of both off-policy and on-policy reinforcement learning algorithms within the context of quantum control. Specifically, the off-policy algorithm, soft actor critic (SAC), although in some task it requires fewer interactions with the environment to reach a specified performance threshold and is capable of learning from scratch, the overall performance is not good enough while pursuing a very high fidelity. This limitation may be attributed to its off-policy nature; although SAC is efficient in exploring the environment, it is less adept at precisely determining specific pulse shapes.
Without demonstrations, PPO also struggles to handle quantum control problems with high-dimensional action spaces, as evidenced by its stagnant average reward during training. Nevertheless, when equipped with demonstrations and a pre-trained policy network, PPO excels. The prior knowledge significantly enhances its practicality, allowing it to converge to high-performance policies. The integration of demonstrations ensures greater training stability and convergence. This feature is crucial, especially in complex environments where prior knowledge can be extremely beneficial. Our findings suggest that PPO, when combined with RLfD, can be extended to more complex quantum control problems with high-dimensional action space.
@@ A.2 PPO algorithm
The algorithm PPO with pre-train for Quantum Control is shown in Algorithm 2.
$H(\cdot|s_{t})$ is the entropy of $\pi_{\theta}(\cdot|s_{t})$, it’s used to maintain randomness of $\pi_{\theta}(\cdot|s_{t})$ to preserve the exploration ability. Entropy here is used to prevent the policy’s variance from being zero. Here $\lambda_{\rm{ent}}$ is the coefficient of entropy that controls the extent of exploration by influencing how much randomness is introduced into the policy.
We finally show the comparison of pulses shapes generated by RL agent and original demonstrations in Fig 5. We can see that the pulse shapes generated by the RL agent are only slightly different from the original demonstrations. These subtle differences indicate that the agent is capable of fine-tuning the pulse shapes to better adapt to the environment. By making minor adjustments, the RL agent can compensate for system imperfections or calibration errors, leading to improved performance in the quantum control task.